OpenStudy (anonymous):

Use the matrix tool to find a solution to the systems of equations. Solve for x. 3x-y-2z=-3 x+2y+z=-7 2x-y+5z=12

OpenStudy (anonymous):

Matrix tool? From what? A graphing calculator? Or something like mathway?

OpenStudy (anonymous):

its an arrangment of columns and rows such as 0 0 0|0 0 0 0|0 0 0 0|0

OpenStudy (anonymous):

I'm well aware of what a matrix is. You're talking about some matrix tool. Any idea what that might be referring to?

OpenStudy (anonymous):

Its just a matrix tool with the capability of dividing subtracting etc. rows

OpenStudy (anonymous):

3 -1 -2 | -3 1 2 1 | -7 2 -1 5 | 12 R1 = R1-R3 1 0 -7 | -15 1 2 1 | -7 2 -1 5 | 12 R2 = R2+2(R3) 1 0 -7 | -15 5 0 11 | 17 2 -1 5 | 12 R1 = R1+11/7(R2) 62/7 0 0 | 11*17/7 -15 ----- ---- so (62/7)x = (11*17)/7-15 x = 41/31 i guess

OpenStudy (anonymous):

Last operation should be R1 = R1+7/11(R2) 46/11 0 0 | -15 + 7/11*17 --- -- so 46/11x = -15 + (7/11)*17 x = -1 That is the correct solution

OpenStudy (anonymous):

I think what they call a matrix tool is not a device or a program but the solution method i used above. They are not hard to solve but takes some time

OpenStudy (anonymous):

thanks i really didnt know ho to solve a three row matrix

OpenStudy (anonymous):

They are not hard to solve. If you want to find x just leave its coefficient alone(make othes 0) by adding and subtracting other rows from any row you chose. After finding something like: a 0 0 | b You can write it as ax + 0y + 0z = b It is equal to ax = b and x = b/a

OpenStudy (anonymous):

In my solution you can see that second element of the first row became 0. Then i tried to make 3rd element 0. But if i add or subtract 2nd or 3rd row the second element would be different than 0. So i prepared the second row first by making the second elemnt of it 0. Then i maked the 3rd element of first row 0

OpenStudy (anonymous):

made (not maked ) :D hahaha

OpenStudy (anonymous):

thanks! it was really helpful.